MISSION STATUS (updated Feb 9, 2002)

Feb 9, 2002, 10:30AM PST:
All systems are working as planned. No problems. From now on, we will reduce the number of announcements in this list. For a detailed description of the passes please check the HESSI Major Events .

Feb 6, 2002, 5:15PM PST:
Soon the solid state recorder will be turned on. Downlink now is 1Mbit per second. The cryocooler is operating at 85% of the power. The particle detector is turned on and working well.

Feb 5, 2002, 10:25 PM PST:
The cryocooler was turned on for the first time

Feb 5, 2002, 5:15 PM PST:
Everything is still hunky dory. The 2nd pass went fine. We are within 5 degrees of the sun. Command uplink and downlink is strong. Next big events: crycooler turn-on (at about 8PM PST) and start spinning up.

Look at the HESSI Major Events page for details about the passes.

Feb 5, 2002, 4:15 PM PST:
Press articles:

Feb 5, 2002, 4:07 PM PST:
From Brian Dennis, HESSI Mission Scientist: I'm pleased to report that HESSI telemetry was picked up at Berkeley and at Wallops on the first orbit and that everything looks nominal. The solar panels have deployed and the spacecraft is pointing close to the solar direction. The orbit is very close to the nominal that Peter Harvey gave in his last email (see below) and the ground stations had no trouble picking up the spacecraft.

Feb 5, 2002, 3:50 PM PST:
The sun is aquired within 15 degrees. Solar panels are deployed. Everything is nominal. Next pass at 4:11 PST.

From Peter Harvey, Project Manager: HESSI was successfully launched today 4:08 PM EST. Based upon launch vehicle data, the orbit is approximately 586.85 x 600.24 km at 38.023 degrees inclination.

The speceraft telemetry was good all the way through launch with no anomalies.

Feb 4, 2002:
Over the weekend, tests of the launch vehicle and spacecraft were successful. We continued charging the battery in preparation for launch. Today, we completed the Launch Readiness Review with all parties giving the 'GO' for launch. Weather is also cooperating.


If all goes as planned, solar images will soon be returned by the HESSI space craft. SOHO image, courtesy of the SOHO project team.


December 2001:

  • OSC replaces rocket 1st stage skirt, motors and fins.
  • New assembly passes vehicle verification tests.
  • Final test of science detectors. All good.
  • Structured warm up.
  • Authorization to proceed at KSC meeting (LVRR).
  • Spacecraft mated with rocket

November 2001:

  • OSC verifies and replaces Fin Actuator System and Thrust Vector Controller.
  • OSC makes rocket software modifications (trajectory control) as a result of X43 failure and wind tunnel data.
  • OSC submits description of software changes to KSC.
  • Interface Verification and Flight Sim 3 testing completed.

October 2001:

  • OSC removes fin control electronics to check for contamination.
  • OSC announces plan for additional wind tunnel testing.
  • OSC carries out additional wind tunnel testing.

September 2001:

  • Spacecraft and detector tests performed.
  • Begin structured warmup using cryocooler.
  • Return to LN2 cooling.
  • Failure of OSC Taurus launch. OSC reviews causes and whether there are implications for Pegasus.
  • Successful Athena launch from Alaska. KSC staff returns to complete review of questions related to X-43.

August 2001:

  • Maintenance cooling with LN2.
  • Regular monitoring of temperatures and pressure.

July 2001:

  • Hessi demated from Pegasus rocket.
  • Structured cooldown using cryocooler.
  • High voltage turned on and detectors checked out. All good.
  • Spacecraft turned vertical, cryocooler turned off and maintenance cooling with LN2.

June 2001:

  • Rocket with Hessi on L-1011 flies to KSC.
  • Post flight testing of rocket and satellite.
  • Continued testing of Berkeley Ground Station.
  • Setup and checkout of Santiago Ground Station.
  • X-43A test flight fails, causing delay of Hessi launch.
  • Hessi returns to VAFB to await launch authorization.

May 2001 (VAFB):

  • Pegasus rocket successfully completes Qualification testing.
  • Performed structured warmup of Hessi detectors.
  • HESSI mated with Pegasus rocket.
  • Fairings installed.
  • Pegasus with Hessi moved to transport trailer.

April 2001 (VAFB):

  • Continuing Attitude Control System simulations and tests.
  • Continuing to practice flight operation procedures.
  • Moved spacecraft to Bldg 1555 at VAFB.
  • Identified and corrected minor anomaly with transponder.

March 2001 (VAFB):

  • Completed setup of Hot Bench at UCB and performed Attitude Control System simulations and tests.
  • Moved spacecraft and GSE to Building 1610 on higher ground at VAFB to avoid power outages and flooding.
  • Completed 4 Day No Command test.
  • Practising flight operations procedures.

February 2001 (VAFB):

  • Attached solar panels and performed 1st Motion Test.
  • Repeated spacecraft and instrument functional tests.
  • Reviewed and tested flight operations procedures.
  • Presented mission status and readiness to PreShipment and Red Team Reviews.

January 2001 (JPL and VAFB):

  • Spacecraft returns to thermal vacuum for testing of cryocooler performance at 50W and amplifier trims. Thermal Vacuum successfully concluded on January 8.
  • Mass Properties Measurements and Spin Balance performed on January 9 and 10.
  • Spacecraft transported to VAFB on January 12.
  • Setup and Safe-to-Mate testing performed January 13-14.
  • Spacecraft functionals and calibrations successfully performed on Jan 15-17.




  • Solar array deployment in thermal vacuum successfully tested.
  • Endurance thermal-vacuum testing including hot and cold dwells and extended thermal balance testing.
  • Learned that spectrometer is colder than expected at planned cryocooler power level. Good news: we can operate with less power.
  • Observed hf noise on two detectors.
  • Made amplifier trims to correct noise problem.


  • Spacecraft transported to JPL on November 1-2.
  • Preliminary setup and testing performed.
  • Solar Arrays attached and deployment tested.
  • Vibration testing of X , Y and Z axes successfully completed.
  • Separation/shock testing successfully completed.


  • Flight battery reconditioned and installed.
  • Spacecraft fully functional and running.
  • Red Team and Pre-Environmental Reviews.


  • Ran and passed the following spacecraft functional tests: command and data handling, inertia adjustment device, thermal, launch vehicle interface, electrical power system.


  • The Spacecraft Electronics Module, Solid State Recorder, Sun Sensors and Magnetometer have been returned from subcontractors and re-installed on the spacecraft. All units checked out OK.
  • Requalified Transponder and RF Switch received and installed on the spacecraft.
  • Safe-to-Mate functional tests run for all components.


  • Solid State Recorder, Transponder, RF Switch, Sun Sensors and Magnetometer all removed from the spacecraft and delivered to the subcontractors for retesting and qualification.
  • The Spacecraft Electronics Module completed testing at Spectrum Astro by the end of the month and checked out OK.


  • New deck completed and installed.
  • Refurbished bus structure completed and aligned.
  • New imager support ring completed and installed.
  • Implementing several plans running in parallel to provide new spectrometer cryocooler.
  • New solar array substrates completed at COI.
  • Requests for Action received from May review team.  Spacecraft Electronics Module shipped back to Spectrum Astro.
  • GPMC authorizes new integration, test and launch plan on condition that cryocooler issue is resolved by September. New launch date set for March 2001.


  • Spacecraft minus imager and spectrometer taken to Bell Labs for thermal test.  From May 2 to May 5 spacecraft undergoes 6 cycles between +45 and -20C.  Spacecraft passes all functional tests except for firing of frangibolts.  Problem is found to be a  broken wire which is fixed.  All other wiring is scrutinized and plans prepared for high pot and other testing.
  • Designed stronger imager support ring.
  • Decided to repair broken solar panels and also contract new panels to have two options and choose the best in future.
  • Honeycomb deck found to be slightly warped.  Began fabrication of new deck.
  • Spectrometer cryocooler found to have deteriorating performance since mishap.
  • Spectrometer cryocooler flight spare starts also exhibiting problems during burn-in tests.
  • Termination/Recovery Review held at UCB on May 15 and 16.
  • HESSI Test Mishap Investigation Board issues their report on May 18.
  • After the thermal test and fix of frangibolt wiring, spacecraft is powered up and running most days.
  • Flight operations team continues preparations,  including simulations of Launch and Early Orbit plus regular passes.


  • Spacecraft carefully inspected and found to be in good shape with the following exceptions:  two damaged solar arrays,  broken imager support ring,  one strained cable.
  • Review meeting between GSFC and UCB held on April 4.  Recommendation is to do more inspection plus a thermal test.
  • Imager (which rests on the broken imager support ring) shipped to PSI so they can fully test and recalibrate.
  • Spectrometer removed and tested for leaks.  None found.


  • Solar arrays integrated and thermal blankets installed.
  • Spacecraft shipped to JPL.
  • Pass EMC testing.
  • Pass random vibration testing.
  • Spacecraft over-vibrated in sine-burst test (20 G's applied instead of 2 G's for 200 milliseconds).
  • Review team gathered to investigate the over-test.
  • Spacecraft shipped back to UCB for inspection and evaluation.



  • Spacecraft Bus delivered to UC Berkeley.
  • All instruments successfully integrated and tested at UCB's Spacecraft Integration Facility.
  • Pass Environmental Reviews.
  • Successful end-to-end testing of telemetry and commands between spacecraft and ground system via the 11 meter antenna.